# Shadow of black holes with consistent thermodynamics

### v1

@@ I Introduction
Black hole solutions represent fundamental spacetime configurations in Einstein's theory of General Relativity (GR). However, the presence of singularities signals the limitations of GR. Indeed, there have been a lot of proposals to modify the metric to resolve singularities in the black hole geometry 1968qtr..conf…87B ; Dymnikova:1992ux ; Hayward:2005gi ; Bambi:2013ufa . For reviews and more references, see Torres:2022twv ; Lan:2023cvz . The resolution of the problem comes most probably from the quantum effects, which might produce such modifications of the metric. In the study of black holes, the associated thermodynamic properties may offer a remarkable window into quantum effects manifesting at the semiclassical level.
@@
One of the important consequences in the asymptotic safety is that the couplings of the theory become functions of an energy scale. In the search for physical implications of this energy scale for the black hole solutions, its identification with some length scale has been considered in the geometry of black holes Bonanno:2000ep ; Bonanno:2006eu ; Reuter:2004nv ; Falls:2010he ; Reuter:2010xb ; Harst:2011zx ; Falls:2012nd ; Koch:2013owa ; Litim:2013gga ; Koch:2014cqa ; Bonanno:2017zen ; Pawlowski:2018swz ; Platania:2020lqb ; Ishibashi:2021kmf ; Ruiz:2021qfp . The identification generally depends on the spacetime coordinates and associated physical quantities. However, there had not been found any fundamental principle that uniquely determines this identification, other than the dimensional reason. On the other hand, when the rotating Kerr black hole is considered, it is found that the first law of black hole thermodynamics is not satisfied straightforwardly Reuter:2010xb . Since the first law of thermodynamics is the fundamental law of energy conservation, it must be valid not only semiclassically but also quantum mechanically. It has been shown that the consistency condition of the first law of the black hole thermodynamics imposes a very general, albeit not so strong, nontrivial constraint on the dependence of the Newton coupling on the black hole parameters and geometry in the quantum theory Chen:2022xjk ; Chen:2023wdg .
@@
A class of black hole observations that is suitable for our purpose is the shadow images cast by black holes Falcke:1999pj ; EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp ; Chen:2022scf . Due to the strong gravitational fields near black holes, the light rays that propagate around the black hole can be highly lensed and orbit the black hole multiple times, as long as their trajectories get close enough to the black hole. The black hole images that appear on the observer's image plane consist of the higher-order images contributed by the highly lensed light rays, on top of the direct emission contributed by the light rays that propagate directly to the observer, without orbiting the black hole Bozza:2010xqn ; Gralla:2019xty ; Johnson:2019ljv . The more highly lensed light rays, as well as their higher-order images, are expected to carry a larger fraction of information about the black hole geometry itself, and are thus less sensitive to the emission profiles under consideration Cunha:2018acu . In particular, as one increases the order of lensed images, the higher-order images quickly converge to a closed contour that we term ``critical curve" [^1] This terminology is adopted from Ref. Gralla:2019drh . on the image plane Gralla:2019drh , which corresponds to the impact parameter of infinitely lensed light rays orbiting on a set of unstable spherical orbits around the black hole. This set of spherical orbits and the shadow critical curve on the images are solely determined by the black hole metric through the photon geodesic equations. Therefore, shadow critical curves carry pure information about the black hole geometry and can be very crucial to probe the black hole geometry itself in a theory-agnostic manner. For nonrotating black holes, it is well known that the shadow critical curve appears as a perfect circle, while a rotating black hole, e.g., the Kerr one, would cast a D-shaped shadow due to the frame-dragging effects Luminet:1979nyg . By properly identifying the observables that describe the size or the shape of the critical curve, one could use them to constrain the parameter space of the model or extract relevant information about the black hole geometry Johannsen:2015hib ; Younsi:2021dxe . This line of research has been widely considered in the literature Wielgus:2021peu ; Vagnozzi:2022moj , for instance, for black holes in modified theories of gravity Zhu:2019ura ; Khodadi:2020gns ; Ghosh:2020spb ; Ghosh:2022kit ; Jusufi:2022loj ; EventHorizonTelescope:2021dqv ; Afrin:2021wlj ; Papnoi:2021rvw ; Pantig:2022ely ; Kuang:2022ojj ; Khodadi:2022pqh ; Meng:2022kjs , black holes with quantum corrections Brahma:2020eos ; Eichhorn:2021iwq ; KumarWalia:2022ddq ; Afrin:2022ztr ; Islam:2022wck ; Zhang:2023okw ; Jiang:2023img ; Cao:2024vtq , horizonless compact objects Dey:2020bgo ; Bouhmadi-Lopez:2021zwt ; Shaikh:2021cvl ; Rahaman:2021web ; Shaikh:2022ivr ; Olmo:2023lil ; Nguyen:2023clb , and several parameterizations of non-Kerr spacetimes Ghasemi-Nodehi:2015raa ; Younsi:2016azx ; Wang:2017hjl ; Shaikh:2019fpu ; EventHorizonTelescope:2020qrl ; Chen:2022lct ; Konoplya:2021slg .
@@
As the first step toward testing the consistency of black hole thermodynamics through black hole observations, in this paper, we consider a class of minimal extensions of the Kerr spacetime. The metric of this model, which was proposed in Ref. Chen:2023wdg , has the form of the usual Kerr one, but the Newton coupling takes the form $G(r,M,a)=G(r^{2}+a^{2},Mr)$, replacing the original Newton coupling constant $G_{0}$. As we will show in Sec. II, this class of minimal extensions ensures the integrability of the first law of black hole thermodynamics and hence the consistency condition. However, depending on the explicit parameterization of the Newton coupling under consideration, the model can still have a huge range of parameter space. In this paper, we take a general parameterization of the Newton coupling $G(r^{2}+a^{2},Mr)$, consider a few representative cases, and perform a thorough analysis of how the shape and size of shadows vary in a wide range of parameter space. Our analysis is twofold: First, we would like to place constraints on the parameter space of the model. This can be achieved by using the available constraint on the observed size of Sgr A* shadow images EventHorizonTelescope:2022xqj , incorporated by the accurate measurements on the mass-to-distance ratio of Sgr A* using surrounding stellar dynamics. Second, a crucial but much more difficult task of this work is to identify possible observational features or implications that can be shared within a sufficiently large (if not all) parameter space of this model. This direction is important in order to draw conclusions regarding the observational implications of the consistency of black hole thermodynamics in general. In particular, if such features or implications exist, they can assist in the attempt to test the consistency conditions of black hole thermodynamics through shadow observations.
@@
The paper is organized as follows. In Sec. II, we review the black hole thermodynamics and the conditions for the black hole metric in order to satisfy the consistency of thermodynamics, from which we introduce the class of minimal extensions of Kerr black holes we will be considering. Then, we introduce our theoretical setup in Sec. III. In Sec. IV, we briefly review the null geodesic equations, the spherical photon orbits, and the derivation of shadow critical curves. Then, in Subsec. IV.2, we introduce two shadow observables, which quantify the apparent size and the distortion of the critical curves, respectively. In Sec. V, we present our results on the black hole shadows in our model. In particular, we identify possible observational implications on the shadow critical curves that could be useful in terms of testing the consistency of black hole thermodynamics through black hole images. Then, in Sec. VI, we discuss the possibility of going beyond the Kerr bound on the black hole spin in our model, and reveal an interesting shadow feature in this case. Finally, we conclude in Sec. VII.
@@ II Consistency of thermodynamics
$ds^{2}=-\frac{\Delta}{\Sigma}\left(dt-a\sin^{2}\theta\,d\varphi\right)^{2}+\frac{\Sigma}{\Delta}\,dr^{2}+\Sigma\,d\theta^{2}+\frac{\sin^{2}\theta}{\Sigma}\left[a\,dt-(r^{2}+a^{2})\,d\varphi\right]^{2}\,$, (1)
@@
It is well known that for the Kerr spacetime, the first law of black hole thermodynamics holds:
@@
where $S$ and $T$ denote the entropy and temperature of the black hole, respectively. The thermodynamic quantities are related as follows:
@@
In this thermodynamic framework, $M$ and $a$ are treated as independent variables. According to the first law, the differential of the entropy can be expressed as:
@@
To ensure internal consistency, the mixed partial derivatives of $S$ must satisfy the integrability condition:
@@
If the Newton coupling is a constant with the horizon radius $r_{h}$ determined from (5), the condition (7) is obviously satisfied. However, when the Newton coupling becomes scale-dependent due to, for instance, quantum effects, i.e., instead of $G=G_{0}$ a constant, when $G=G(r,M,a)$, this condition imposes nontrivial constraint on $G$. To see this, let us first note that the temperature is computed via the surface gravity $\kappa$, giving:
@@
In addition, on the horizon, the condition (5) implicitly defines $r_{h}=r_{h}(M,a)$. It is then helpful to express all relevant derivatives, $\partial_{M},\partial_{a},\partial_{r_{h}}$, and use the chain rule to relate them. In the analysis in Chen:2023wdg , it is shown that the integrability condition (7) for the first law leads to
@@
This constraint must be satisfied in order for the running Newton coupling to fulfill the thermodynamic consistency. A general solution to this condition is:
@@
$S=\int\frac{dA}{4\left(G+\partial_{M}G\right)}\,$, (11)
@@ III Setup
Based on the discussion in the previous section, we consider the rotating black hole spacetime described by the line element (1), with the original constant Newton coupling $G_{0}$ replaced by a running Newton coupling $G(r,M,a)$. Depending on the explicit form of $G(r,M,a)$, there may be one or multiple horizons, which can be determined by solving the equation $\Delta(r_{h})=0$. The outer event horizon $r_{h}$, which is relevant in the context of black hole thermodynamics, corresponds to the largest real solution of the equation $\Delta(r_{h})=0$.
For later convenience, we define the following dimensionless parameter
@@
The extremal configuration exists only if there is a real solution to Eq. (13) [and (5)]. Such a configuration can be obtained by solving Eqs. (5) and (13). One gets
@@
In the unit $c=\hbar=1$ such that $\textrm{Time}]=[\textrm{Length}]=1/[\textrm{Mass}]$, both the terms $G_{0}/(r^{2}+a^{2})$ and $Mr$ are dimensionless. A convenient way of parameterizing $G(r,M,a)$ is to introduce the variable
@@
    One can see that when $p+l>0$ and for a given $a_{*}$, the contributions from the running of $G$ through the dependence on $z$ will be suppressed by the black hole mass $M$, i.e., $z\propto 1/M^{2p+2l}$. For supermassive black holes, the contributions from the running of $G$ can then be unmeasurably small Glampedakis:2021oie . However, if (l,p)=(1,1)(l,p)=(1,-1), or in general p=lp=-l, such a mass suppression does not occur. In this case, all the coupling constants that appear in $G(z)$ can be directly constrained through black hole observations, irrespective of the mass of the black holes being observed.
@@
    where $\tilde{\omega}>0$ is a coupling constant. At spatial infinity $r\rightarrow\infty$, we have $z\rightarrow 0$ and $G(z)\rightarrow G_{0}$. While near the origin $r=0$, we have $z\rightarrow\infty$ and $G(z)$ approaches zero, hence the singularity could be resolved. In this work, we will consider $(l,p)=(1,4)$ as an example inspired by this type of nonsingular black hole models.
@@ IV Null geodesic and shadow observables
The main aim of this paper is to investigate the shadows cast by the black holes that satisfy the thermodynamic consistency. In what follows, we will focus on the black hole metric given by Eq. (1), with a running Newton coupling (15). In particular, we will study the shadow critical curve on the image plane as seen by a distant observer. Essentially, the critical curve is defined by the impact parameters of the unstable spherical photon orbits around the black hole. These orbits are unstable in the sense that if one slightly perturbs the photons moving on them, these photons would either escape away from the black hole or fall into the horizon. In particular, they represent the asymptotic limit of orbits on which light rays are infinitely lensed. Most importantly, they are solely determined by the black hole spacetime itself and do not depend on the emission models or other astrophysical environments around the black hole. In this section, we will first review the null geodesic equations for the spacetime (1) and derive the equations that determine the shadow critical curves (see also Ref. Tsukamoto:2017fxq ). Then, we will introduce the observables that quantify the size and shape of the shadow critical curves.

### v2

# Shadow of rotating black holes with consistent thermodynamics

@@ I Introduction
Black hole solutions represent fundamental spacetime configurations in Einstein's theory of General Relativity (GR). However, the presence of singularities signals the limitations of GR. Indeed, there have been a lot of proposals to modify the metric to resolve singularities in the black hole geometry 1968qtr..conf…87B ; Dymnikova:1992ux ; Hayward:2005gi ; Bambi:2013ufa . For reviews and more references, see Refs. Torres:2022twv ; Lan:2023cvz . The resolution of the problem comes most probably from the quantum effects, which might produce such modifications of the metric. In the study of black holes, the associated thermodynamic properties may offer a remarkable window into quantum effects manifesting at the semiclassical level.
@@
One of the important consequences in the asymptotic safety is that the couplings of the theory become functions of an energy scale. In the search for physical implications of this energy scale for the black hole solutions, its identification with some length scale has been considered in the geometry of black holes Bonanno:2000ep ; Bonanno:2006eu ; Reuter:2004nv ; Falls:2010he ; Reuter:2010xb ; Harst:2011zx ; Falls:2012nd ; Koch:2013owa ; Litim:2013gga ; Koch:2014cqa ; Bonanno:2017zen ; Pawlowski:2018swz ; Platania:2020lqb ; Ishibashi:2021kmf ; Ruiz:2021qfp ; Lu:2024ppa . The identification generally depends on the spacetime coordinates and associated physical quantities. However, any fundamental principle that uniquely determines this identification had not been found, other than the dimensional reason. On the other hand, when the rotating Kerr black hole is considered, it is found that the first law of black hole thermodynamics is not satisfied straightforwardly Reuter:2010xb . Since the first law of thermodynamics is the fundamental law of energy conservation, it must be valid not only semiclassically but also quantum mechanically. It has been shown that the consistency condition of the first law of the black hole thermodynamics imposes a very general, albeit not so strong, nontrivial constraint on the dependence of the Newton coupling on the black hole parameters and geometry in the quantum theory Chen:2022xjk ; Chen:2023wdg .
@@
A class of black hole observations that is suitable for our purpose is the shadow images cast by black holes Falcke:1999pj ; EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp ; Chen:2022scf . Because of the strong gravitational fields near black holes, the light rays that propagate around the black hole can be highly lensed and orbit the black hole multiple times, as long as their trajectories get close enough to the black hole. The black hole images that appear on the observer's image plane consist of the higher-order images contributed by the highly lensed light rays, on top of the direct emission contributed by the light rays that propagate directly to the observer, without orbiting the black hole Bozza:2010xqn ; Gralla:2019xty ; Johnson:2019ljv . The more highly lensed light rays, as well as their higher-order images, are expected to carry a larger fraction of information about the black hole geometry itself, and are thus less sensitive to the emission profiles under consideration Cunha:2018acu . In particular, as one increases the order of lensed images, the higher-order images quickly converge to a closed contour that we term ``critical curve" [^1] This terminology is adopted from Ref. Gralla:2019drh . on the image plane Gralla:2019drh , which corresponds to the impact parameter of infinitely lensed light rays orbiting on a set of unstable spherical orbits around the black hole. This set of spherical orbits and the shadow critical curve on the images are solely determined by the black hole metric through the photon geodesic equations. Therefore, shadow critical curves carry pure information about the black hole geometry and can be very crucial to probe the black hole geometry itself in a theory-agnostic manner. For nonrotating black holes, it is well known that the shadow critical curve appears as a perfect circle, while a rotating black hole, e.g., the Kerr one, would cast a D-shaped shadow due to the frame-dragging effects Luminet:1979nyg . By properly identifying the observables that describe the size or the shape of the critical curve, one could use them to constrain the parameter space of the model or extract relevant information about the black hole geometry Johannsen:2015hib ; Younsi:2021dxe . This line of research has been widely considered in the literature Wielgus:2021peu ; Vagnozzi:2022moj , for instance, for black holes in modified theories of gravity Zhu:2019ura ; Khodadi:2020gns ; Ghosh:2020spb ; Ghosh:2022kit ; Jusufi:2022loj ; EventHorizonTelescope:2021dqv ; Afrin:2021wlj ; Papnoi:2021rvw ; Pantig:2022ely ; Kuang:2022ojj ; Khodadi:2022pqh ; Meng:2022kjs ; Khodadi:2024ubi , black holes with quantum corrections Brahma:2020eos ; Eichhorn:2021iwq ; KumarWalia:2022ddq ; Afrin:2022ztr ; Islam:2022wck ; Zhang:2023okw ; Jiang:2023img ; Cao:2024vtq ; Contreras:2019cmf ; Koch:2025gaw , horizonless compact objects Dey:2020bgo ; Bouhmadi-Lopez:2021zwt ; Shaikh:2021cvl ; Rahaman:2021web ; Shaikh:2022ivr ; Olmo:2023lil ; Nguyen:2023clb , and several parametrizations of non-Kerr spacetimes Ghasemi-Nodehi:2015raa ; Younsi:2016azx ; Wang:2017hjl ; Shaikh:2019fpu ; EventHorizonTelescope:2020qrl ; Chen:2022lct ; Konoplya:2021slg .
@@
As the first step toward testing the consistency of black hole thermodynamics through black hole observations, in this paper, we consider a class of minimal extensions of the Kerr spacetime. The metric of this model, which was proposed in Ref. Chen:2023wdg , has the form of the usual Kerr one, but the Newton coupling takes the form $G(r,M,a)=G(r^{2}+a^{2},Mr)$, replacing the original Newton coupling constant $G_{0}$. As we will show in Sec. II, this class of minimal extensions ensures the integrability of the first law of black hole thermodynamics and hence the consistency condition. However, depending on the explicit parametrization of the Newton coupling under consideration, the model can still have a huge range of parameter space. In this paper, we take a general parametrization of the Newton coupling $G(r^{2}+a^{2},Mr)$, consider a few representative cases, and perform a thorough analysis of how the shape and size of shadows vary in a wide range of parameter space. Our analysis is twofold: First, we would like to place constraints on the parameter space of the model. This can be achieved by using the available constraint on the observed size of Sgr A* shadow images EventHorizonTelescope:2022xqj , incorporated by the accurate measurements on the mass-to-distance ratio of Sgr A* using surrounding stellar dynamics. Second, a crucial but much more difficult task of this work is to identify possible observational features or implications that can be shared within a sufficiently large (if not all) parameter space of this model. This direction is important in order to draw conclusions regarding the observational implications of the consistency of black hole thermodynamics in general. In particular, if such features or implications exist, they can assist in the attempt to test the consistency conditions of black hole thermodynamics through shadow observations.
@@
The paper is organized as follows. In Sec. II, we review the black hole thermodynamics and the conditions for the black hole metric in order to satisfy the consistency of thermodynamics, from which we introduce the class of minimal extensions of Kerr black holes we will be considering. Then, we introduce our theoretical setup in Sec. III. In Sec. IV, we briefly review the null geodesic equations, the spherical photon orbits, and the derivation of shadow critical curves. Then, in Sec. IV.2, we introduce two shadow observables, which quantify the apparent size and the distortion of the critical curves, respectively. In Sec. V, we present our results on the black hole shadows in our model. In particular, we identify possible observational implications on the shadow critical curves that could be useful in terms of testing the consistency of black hole thermodynamics through black hole images. Then, in Sec. VI, we discuss the possibility of going beyond the Kerr bound on the black hole spin in our model, and reveal an interesting shadow feature in this case. Finally, we conclude in Sec. VII.
@@ II Consistency of thermodynamics
$ds^{2}=$ $-\frac{\Delta}{\Sigma}\left(dt-a\sin^{2}\theta\,d\varphi\right)^{2}+\frac{\Sigma}{\Delta}\,dr^{2}$
$+\Sigma\,d\theta^{2}+\frac{\sin^{2}\theta}{\Sigma}\left[a\,dt-(r^{2}+a^{2})\,d\varphi\right]^{2}\,$, (1)
@@
It is well known that for the Kerr spacetime the first law of black hole thermodynamics holds:
@@
where $S$ and $T$ denote the entropy and temperature of the black hole, respectively. The thermodynamic quantities are related as
@@
In this thermodynamic framework, $M$ and $a$ are treated as independent variables. According to the first law, the differential of the entropy can be expressed as
@@
To ensure internal consistency, the mixed partial derivatives of $S$ must satisfy the integrability condition,
@@
If the Newton coupling is a constant with the horizon radius $r_{h}$ determined from (5), the condition (7) is obviously satisfied. However, when the Newton coupling becomes scale dependent due to, for instance, quantum effects, i.e., instead of $G=G_{0}$ a constant, when $G=G(r,M,a)$, this condition imposes nontrivial constraint on $G$. To see this, let us first note that the temperature is computed via the surface gravity $\kappa$, giving
@@
In addition, on the horizon, the condition (5) implicitly defines [^2] For an asymptotically flat black hole with the ADM mass given by $M$ and the Newton coupling going to the finite constant $G_{0}$ in the asymptotic region, the energy $E$ and the horizon angular velocity $\Omega\equiv|g_{t\varphi}|/g_{\varphi\varphi}$ on the horizon $r_{h}$ are still given by Eq. (4). $r_{h}=r_{h}(M,a)$. It is then helpful to express all relevant derivatives, $\partial_{M},\partial_{a},\partial_{r_{h}}$, and use the chain rule to relate them. In the analysis in Ref. Chen:2023wdg , it is shown that the integrability condition (7) for the first law leads to
@@
This constraint must be satisfied in order for the running Newton coupling to fulfill the thermodynamic consistency. A general solution to this condition is
@@
$S=\int\frac{dA}{4\left(G+M\partial_{M}G\right)}\,$, (11)
@@ III Setup
Based on the discussion in the previous section, we consider the rotating black hole spacetime described by the line element (II), with the original constant Newton coupling $G_{0}$ replaced by a running Newton coupling $G(r,M,a)$. Depending on the explicit form of $G(r,M,a)$, there may be one or multiple horizons, which can be determined by solving the equation $\Delta(r_{h})=0$. The outer event horizon $r_{h}$, which is relevant in the context of black hole thermodynamics, corresponds to the largest real solution of the equation $\Delta(r_{h})=0$.
For later convenience, we define the dimensionless parameter
@@
The extremal configuration exists only if there is a real solution to Eq. (13) [and Eq. (5)]. Such a configuration can be obtained by solving Eqs. (5) and (13). One gets
@@
In the natural unit $c=\hbar=1$ that we use, $\textrm{Time}]=[\textrm{Length}]=1/[\textrm{Mass}]$ and both the terms $G_{0}/(r^{2}+a^{2})$ and $Mr$ are dimensionless. However, this unit obscures where the quantum effects are. It is not difficult to recover the dependence on $\hbar$ and $c$, and we find
$\frac{G_{0}}{r^{2}+a^{2}}\Rightarrow\frac{\hbar\,G_{0}}{(r^{2}+a^{2}/c^{2})c^{3}}\,;\qquad\frac{1}{Mr}\Rightarrow\frac{\hbar}{Mrc}\,.$ (16)
Thus we find that the scale dependence indeed disappears if we take the classical limit $\hbar\to 0$. That being said, let us use the natural unit again for simplicity of presentation.
A convenient way of parametrizing $G(r,M,a)$ is to introduce the variable
@@
    One can see that when $p+l>0$ and for a given $a_{*}$, the contributions from the running of $G$ through the dependence on $z$ will be suppressed by the black hole mass $M$, i.e., $z\propto 1/M^{2p+2l}$. After restoring $\hbar$ in the expression, it becomes clear that this corresponds to the suppression of quantum effects by the black hole mass at the horizon scale, i.e., $z\propto\left(M_{p}/M\right)^{2p+2l}$, with $M_{p}$ the Planck mass. For supermassive black holes, the contributions from the running of $G$ can then be unmeasurably small Glampedakis:2021oie . However, if (l,p)=(1,1)(l,p)=(1,-1), or in general p=lp=-l, such a mass suppression does not occur. [^4] Actually, this is precisely the combination that $\hbar$ drops out. However, as emphasized in the Introduction, we can consider the scale dependence in a broader context with thermodynamic consistency. In this case, all the coupling constants that appear in $G(z)$ can be directly constrained through black hole observations, irrespective of the mass of the black holes being observed.
@@
    where $\tilde{\omega}>0$ is a coupling constant. At spatial infinity $r\rightarrow\infty$, we have $z\rightarrow 0$ and $G(z)\rightarrow G_{0}$. On the other hand, near the origin $r=0$, we have $z\rightarrow\infty$ and $G(z)$ approaches zero, hence the singularity could be resolved. In this work, we will consider $(l,p)=(1,4)$ as an example inspired by this type of nonsingular black hole models.
@@ IV Null geodesic and shadow observables
The main aim of this paper is to investigate the shadows cast by the black holes that satisfy the thermodynamic consistency. In what follows, we will focus on the black hole metric given by Eq. (II), with a running Newton coupling (15). In particular, we will study the shadow critical curve on the image plane as seen by a distant observer. Essentially, the critical curve is defined by the impact parameters of the unstable spherical photon orbits around the black hole. These orbits are unstable in the sense that if one slightly perturbs the photons moving on them, these photons would either escape away from the black hole or fall into the horizon. In particular, they represent the asymptotic limit of orbits on which light rays are infinitely lensed. Most importantly, they are solely determined by the black hole spacetime itself and do not depend on the emission models or other astrophysical environments around the black hole. In this section, we will first review the null geodesic equations for the spacetime (II) and derive the equations that determine the shadow critical curves (see also Ref. Tsukamoto:2017fxq ). Then, we will introduce the observables that quantify the size and shape of the shadow critical curves.